OpenStudy (anonymous):

write as the sum and/or difference of logarithms. express powers as a factor. log4(√x/64) answer choices are a) log4^(x-3 b) -3log4^x c) 12-(1/2)log4^x d) (1/2)log4^(x-3)

OpenStudy (anonymous):

Since it is a quotient, we'll express as a difference: log4(√x/64) = log4(√x) - log4(64) log4(√x/64) = log4 (x)/2 - log4 (4^3) log4(√x/64) = log4 (x)/2 - 3

OpenStudy (anonymous):

You have a few of these questions posted. Just to let you know the rules of logs, which will answer all these questions \[\log A +\log B = \log AB\] \[\log A - \log B = \log \frac{A}{B}\] \[\log A^C = C \log A\] \[\log_x A = n \textrm{ means that } x^n=A \]